OpenStudy (anonymous):

Use Euler's method with step size 0.2 to estimate y(1.4), where y(x) is the solution of the initial-value problem y'=x-xy, y(1)=0 \[y_1=0.2\] \[y_2=0.392\] \[y_3=0.56224\] \[y_4=0.70232\]

OpenStudy (anonymous):

Do I continue until the numbers stop changing or can I stop at y(1.4)?

OpenStudy (anonymous):

It's only asking you to approximate y(1.4) so you don't need to go any further.

OpenStudy (anonymous):

Thank you sir.

OpenStudy (anonymous):

What x value did you begin at? 1. At a step size of .2, that means I'll need to do two steps to get to 1.4

OpenStudy (anonymous):

Correct. My x0=1. I don't know why I added a 4th line. Sry

OpenStudy (anonymous):

You actually don't even need the third line.

OpenStudy (anonymous):

Yes I do. \[0+0.2(1-0)=0.2\] \[0.2+0.2(1.2-(0.2)(1.2))=0.392\] \[0.392+0.2(1.4-(1.4)(0.392))=0.56224\]

OpenStudy (anonymous):

I'm pretty confident that you don't. You KNOW the y value when x is 1. That's given. Your step size is .2, and each time you do your fancy-pants calculation, it gives you an approximation for what y is when x is .2 higher than before.